# AdaNCA: Neural Cellular Automata As Adaptors For More Robust Vision Transformer

### v4

@@ 3.2 AdaNCA architecture
Figure 3: Overview of AdaNCA architecture.
The architecture of AdaNCA is shown in Figure 3. It shares a similar update scheme with the standard NCA as described in Equation 7, but is more computationally efficient due to the proposed Dynamic Interaction stage. All convolutional kernels for token interaction are trainable. In the following paragraphs, we first present the design of the Dynamic Interaction stage and then introduce a way for more efficient token interaction using multi-scale Dynamic Interaction.
@@ 3.2.1 Dynamic Interaction
The Interaction stage in the original NCA performs channel-wise concatenation of the results from different depth-wise convolutions. Instead, our Dynamic Interaction computes a weighted sum of those results. Specifically, a weight computation module $\mathcal{W}_{\mathcal{I}}$ takes the token map as input and outputs per-token scalar weights $\mathbf{W}_{\mathcal{I}m}\in\mathbb{R}^{H\times W\times 1}$. We modify Equation 3 to Equation 8:
$\mathbf{S}_{D\mathcal{I}}=\sum(\mathbf{S}*\mathcal{W}_{\mathcal{I}})(\mathbf{S}\circledast[\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{\mathcal{M}}])=\sum(\mathbf{S}*\mathcal{W}_{\mathcal{I}})\mathbf{S}_{m}=\sum_{m=1}^{\mathcal{M}}\sum_{c=1}^{C}\mathbf{W}_{\mathcal{I}m}\odot\mathbf{S}_{mc},$ (8)
where $\mathbf{S}_{mc}\in\mathbb{R}^{H\times W\times 1},\mathbf{S}_{D\mathcal{I}},\mathbf{S}_{m}\in\mathbb{R}^{H\times W\times C}$. We instantiate the weight computation module $\mathcal{W}_{\mathcal{I}}$ using a two-layer convolutional network. The first layer transforms the dimensionality from $C$ to $\mathcal{M}$, and the second layer computes the actual weights, producing $\mathcal{M}$ scalars for each token. Both layers use $3\times 3$ convolutions to factor in information from both the token and its neighbors. To stabilize training, we add a batch normalization between the two convolutions. Our design of the weight computation network coincides with the one in [22], although our focus is on extracting various information from the same neighborhoods rather than aggregating data from different neighborhoods.
@@ 3.2.2 Multi-scale Dynamic Interaction
Inspired by [47] that uses multi-scale token Interaction to facilitate long-range token communication, we propose multi-scale Dynamic Interaction. Concretely, all convolutions in Equation 8 now have one more degree of freedom in dilation. Dilation $s$ represents the current operating scale being $s$, and $s\in\{1,2,...,\mathcal{S}\}$. Hence, the original Dynamic Interaction is a special case where $\mathcal{S}=1$. To increase the feature expressivity, we perform a weighted sum on the outputs of all scales, where the per-token weights $\mathbf{W}_{Ms}$ are generated by a network $\mathcal{W}_{M}$ as described in Equation 9.
$\mathbf{S}_{MD\mathcal{I}}=\sum(\mathbf{S}*\mathcal{W}_{M})\mathbf{S}_{D\mathcal{I}}=\sum(\mathbf{S}*\mathcal{W}_{M})\mathbf{S}_{D\mathcal{I}}=\sum_{s=1}^{\mathcal{S}}\sum_{c=1}^{C}\mathbf{W}_{Ms}\odot\mathbf{S}_{D\mathcal{I}sc},$ (9)
where $\mathbf{W}_{Ms},\mathbf{S}_{D\mathcal{I}sc}\in\mathbb{R}^{H\times W\times 1},\mathbf{S}_{MD\mathcal{I}},\mathbf{S}_{D\mathcal{I}s}\in\mathbb{R}^{H\times W\times C}$. The weight computation network $\mathcal{W}_{M}$ shares the same structure with $\mathcal{W}_{\mathcal{I}}$ in the Dynamic Interaction.
@@ 3.3 Insert positions of AdaNCA
Given a ViT and an AdaNCA, we need to determine where to insert AdaNCA to maximize the robustness improvements. To this end, we first establish the correlation between the placement of AdaNCA and the robustness enhancement it brings. Motivated by the fact that the network redundancy contributes to the model robustness [27], we hypothesize that the effect of AdaNCA should correlate to the layer redundancy corresponding to the insert position. To quantify the redundancy, we propose the Set Cohesion Index $\kappa$. Given a trained model with $L$ layers and two layer indices $i,j\in\{1,2,...,L\}$ where $i<j$, $\kappa(i,j)$ is defined in Equation 10.
@@
$\mathbbm{1}$ stands for the indicator function. $Sim(m,n)$ is the function for quantifying the output similarity between layer $m$ and $n$. We choose Centered Kernel Alignment (CKA) [34], a common metric for measuring layer similarities inside or between neural networks [24]. A higher $\kappa$ stands for a more cohesive layer set defined by layers from $i$ to $j$. Inserting AdaNCA after layer $i$ would partition the network into 2 layer sets, and we can compute the sum of $\kappa$ of the layers before and after $i$, i.e. $\mathcal{K}(i)=\kappa(1,i)+\kappa(i+1,L)$, serving as a quantification of the network redundancy corresponding to position $i$. We assume that AdaNCA will not change the layer similarity structure since it is too small compared to a single layer in all ViTs.
On the other hand, we quantify the robustness improvement brought by AdaNCA using the relative increase in the attack failure rate of the AdaNCA-inserted models and the corresponding baseline. Specifically, if a model can achieve $\alpha$ clean test accuracy as well as $\alpha^{\prime}$ accuracy under adversarial attacks, the attack failure rate is $\beta=\frac{\alpha^{\prime}}{\alpha}$. The robustness improvement $\gamma$ is then defined as $\gamma=\frac{\beta_{AdaNCA{}}-\beta_{base}}{\beta_{base}}$. In our experiments, we find that $\gamma$ is significantly correlated with the network redundancy $\mathcal{K}$ ($r=0.6938,p<0.001$). We refer readers to Appendix A for details of the experiments. The results validate our hypothesis and indicate that AdaNCA should be inserted into the position that can maximize network redundancy. We develop a dynamic programming algorithm to find those positions and refer readers to Appendix A.1 for the details.
@@ 4 Experiments
We choose 4 ViT models as the baseline: Swin-base (Swin-B) [37], FAN-Base-Hybrid (FAN-B) [77], RVT-Base-Plus (RVT-B) [41], and ConViT-Base (ConViT-B) [16]. Our choices include a hierarchical model (Swin), a convolution-attention hybrid model (FAN), and two regular models in which all layers share the same structure (RVT and ConViT). RVT is specifically built to be a robust model while ConViT is not. All 4 models are already equipped with different kinds of local structures. Moreover, we choose two SOTA models in terms of robustness against out-of-distribution (OOD) data, TAPADL-RVT and TAPADL-FAN [22] for comparison. Note that the SOTA method involves training with an additional loss (ADL) while we do not incorporate it in our training, since our focus is on the effect of architectural changes. We follow the training scheme for each model, respectively, to train the AdaNCA-equipped model from scratch on ImageNet1K. We conduct the analysis presented in Section 3.3 to decide the optimal position to insert multiple AdaNCA modules, in which the ImageNet1K pre-trained weights for analysis are obtained from the PyTorch Image Models library [68]. To balance between the computational cost and robustness improvement, we limit the number of AdaNCA to 2 or 3, depending on the model architecture. The recurrent time step of all AdaNCA is chosen from $\{[2,2],[2,4],[3,5]\}$, and we follow the design principle of ViT, which is to put more computation in the middle or high layers [37]. We fix the random steps during test to a single integer chosen from the range to get precise results and ensure non-stochasticity during adversarial attacks. All the activation functions used in the MLP in AdaNCA are GELU [28], and the input, hidden, and output dimensionalities of the MLP are all the same. We refer readers to Appendix C.6 for the details of the training and insert scheme of AdaNCA. All of our experiments are performed on 4 Nvidia A100 GPUs.
4.1 Results on Image Classification
Table 1: The performance of AdaNCA-enhanced ViTs and their corresponding baselines. AdaNCA consistently improves the clean accuracy as well as the robustness to adversarial and out-of-distribution inputs of the baseline models. Note that our models also outperform the larger baselines ( sign), indicating that the performance improvement does not merely originate from the increase in the number of parameters or FLOPs. The TAPADL method [22] can lead to more vulnerable models compared to the baselines (green numbers). : the test is conducted on models pre-trained on ImageNet22K [41], see Appendix C.9.
@@ 4.1 Results on Image Classification
(M) (G) Clean Acc. PGD CW APGD-DLR APGD-CE IM-A IM-C ($\downarrow$) IM-R IM-SK
@@
Table 2: Comparisons of corruption error on each corruption type of ImageNet-C. AdaNCA consistently improves the performance of the baseline model in all categories (Swin-B), and can achieve better results in most categories compared to the SOTA method (TAPADL-RVT).
@@ 4.2 Layer similarity structure
Our key assumption in Section 3.3 is that AdaNCA will not change the layer similarity structure due to its small size, and that is why we use pre-trained networks to conduct the analysis. Here, we examine the pair-wise layer similarities in Swin-B [37] and Swin-B-AdaNCA in Figure 5.
Figure 4: Pair-wise layer similarities. Layer sets are marked in red boxes. Swin-B-AdanCA has a clearer stage partition.
Figure 5: Ablation on the (a) scales and (b) number of kernels used in our multi-scale Dynamic Interaction. We choose $\mathcal{S}=2$, $\mathcal{M}=4$ to balance between the clean accuracy and robustness.
@@ 4.3 Ablation studies
We conduct ablation studies on ImageNet100, a 100-class subset of ImageNet1K. Previous studies [7, 15, 71, 73] have shown that ImageNet100 serves as a representative subset of ImageNet1K. Hence, we can obtain representative results for the self-evaluation of the model while efficiently using the computational resources. All the ablation experiments are based on a Swin-tiny [37] model. We insert it after the fourth layer to obtain the best robustness improvement, according to our analysis in Section 3.3 and Appendix A. First, we ablate on two hyperparameters, the number of convolutional kernels used in the Dynamic Interaction stage ($\mathcal{M}$) and the maximum scale ($\mathcal{S}$) used in the multi-scale Dynamic Interaction stage. The Clean Accuracy and Attack Failure Rate are shown in Figure 5. The attack failure rate is quantified using the same method as in Section 3.3 and Appendix A. Multi-scale interaction can contribute to the performance while overly large scale can lose local information. Increasing the number of kernels benefits the performance while too many kernels undermine the robustness. According to the results, we choose $\mathcal{S}=2$, $\mathcal{M}=4$. We then perform ablations on several design choices:
@@
  • Dynamic interaction (DynIn). Ablate the Dynamic Interaction that the interaction results are simply summed together. The number of kernels remains the same.
Results are shown in Table 3. Our design choices contribute to model performance and robustness.
Table 3: Ablation studies on design choices. Each of the AdaNCA design choices contributes to improved performance and robustness. Baseline is a Swin-tiny [37] model trained on ImageNet-100.
Exp. Type Recur StocU RandS DynIn Params (M) FLOPs (G) Accuracy Attack Failure Rate
Baseline 27.69 4.5 86.56 12.29
@@ 4.4 Noise sensitivity examination
A drawback of adversarially robust models is their increased sensitivity to noise in specific frequency bands [59]. For instance, while adversarially trained models are robust against adversarial attacks, they can be sensitive to noise from a larger frequency-band range compared to standard models [59]. Here, we use the method and data from [59] to examine the noise sensitivity of AdaNCA-enhanced ViTs. Specifically, we evaluate the classification performance on a set of images that are mixed with noise of varying magnitudes and frequencies.
Figure 6: Classification accuracy of humans and ViTs on noisy images. AdaNCA improves the accuracy on images with certain types of noise (yellow boxes), indicating that it makes ViTs less sensitive to them. Quantitative: Appendix C.14.
Higher classification accuracy on a specific noise type indicates that the model is less sensitive to that noise. Figure 6 presents the results, including human data sourced from [59]. Our findings demonstrate that AdaNCA enhances ViTs by reducing the sensitivity to noise with certain frequency components, equipping ViTs with more human-like noise-resilient abilities. Crucially, AdaNCA improves the model robustness differently than adversarial training since the AdaNCA-enhanced ViTs do not exhibit increased sensitivity to the noise. We quantitatively validate the conclusion and refer readers to Appendix C.14 for the details.

### v5

@@ 3.2 AdaNCA architecture
Figure 3: Overview of AdaNCA architecture. Instead of concatenating the interaction results generated by the depth-wise convolutions, our Dynamic Interaction conducts a point-wise weighted sum on them to improve the efficiency and enhance the performance. The weights are obtained based on the token states so that each token can dynamically adjust, according to the inputs, the interaction strategy. The Multi-scale Dynamic Interaction aggregates the results from Dynamic Interaction , where the convolutions have different dilation rates. Then, to finish one step of evolution, the output is fed into the Update stage .
The architecture of AdaNCA is shown in Figure 3. It shares a similar update scheme with the standard NCA, as described in Equation 7; but, it is more computationally efficient due to the proposed Dynamic Interaction stage. All convolutional kernels for token interaction are trainable. In the following paragraphs, we first present the design of the Dynamic Interaction stage and then introduce a way for more efficient token interaction by using multi-scale Dynamic Interaction.
@@ 3.2.1 Dynamic Interaction
The Interaction stage in the original NCA performs a channel-wise concatenation of the $\mathcal{M}$ output tensors from different depth-wise convolutions. Whereas, our Dynamic Interaction computes a weighted sum of those results. Specifically, a weight computation network $\mathcal{W}_{\mathcal{I}}$ takes the token map $\mathbf{S}\in\mathbb{R}^{H\times W\times C}$ as input and outputs per-token scalar weights $\mathbf{W}_{\mathcal{I}m}\in\mathbb{R}^{H\times W\times 1}$ for each of the $\mathcal{M}$ kernels. We modify Equation 3 to Equation 8:
$\mathbf{S}_{D\mathcal{I}}=\sum(\mathbf{S}*\mathcal{W}_{\mathcal{I}})(\mathbf{S}\circledast[\mathcal{C}_{1},\mathcal{C}_{2},...,\mathcal{C}_{\mathcal{M}}])=\sum(\mathbf{S}*\mathcal{W}_{\mathcal{I}})\mathbf{S}_{m}=\sum_{m=1}^{\mathcal{M}}\left(\sum_{c=1}^{C}\mathbf{W}_{\mathcal{I}m}\odot\mathbf{S}_{mc}\right),$ (8)
[…]
@@ 3.2.2 Multi-scale Dynamic Interaction
Inspired by [48], which uses multi-scale token Interaction to facilitate long-range token communication, we propose multi-scale Dynamic Interaction. Concretely, all convolutions in Equation 8 now have one more degree of freedom in dilation. Dilation $s$ represents the current operating scale being $s$, and $s\in\{1,2,...,\mathcal{S}\}$. Hence, the original Dynamic Interaction is a special case where $\mathcal{S}=1$. To increase the feature expressivity, we perform a weighted sum on the outputs of all scales, where the per-token weights $\mathbf{W}_{Ms}$ are generated by a network $\mathcal{W}_{M}$ as described in Equation 9.
$\mathbf{S}_{MD\mathcal{I}}=\sum(\mathbf{S}*\mathcal{W}_{M})\mathbf{S}_{D\mathcal{I}}=\sum(\mathbf{S}*\mathcal{W}_{M})\mathbf{S}_{D\mathcal{I}}=\sum_{s=1}^{\mathcal{S}}\left(\sum_{c=1}^{C}\mathbf{W}_{Ms}\odot\mathbf{S}_{D\mathcal{I}sc}\right),$ (9)
where $\mathbf{W}_{Ms},\mathbf{S}_{D\mathcal{I}sc}\in\mathbb{R}^{H\times W\times 1},\mathbf{S}_{MD\mathcal{I}},\mathbf{S}_{D\mathcal{I}s}\in\mathbb{R}^{H\times W\times C}$. The $\mathcal{W}_{\mathcal{I}}$ in Equation 8 is shared across all scales. The weight computation network $\mathcal{W}_{M}$ mirrors that of $\mathcal{W}_{\mathcal{I}}$.
@@ 3.3 Insert positions of AdaNCA
Given a ViT and an AdaNCA, to maximize the robustness improvements, we need to determine where to insert AdaNCA. To this end, we first establish the correlation between the placement of AdaNCA and the robustness enhancement it brings. Motivated by the fact that the network redundancy contributes to the model robustness [28], we hypothesize that the effect of AdaNCA should correlate to the layer redundancy corresponding to the insert position. To quantify the redundancy, we propose the Set Cohesion Index $\kappa$. Given a trained model with $L$ layers and two layer indices $i,j\in\{1,2,...,L\}$ where $i<j$, $\kappa(i,j)$ is defined in Equation 10.
@@
$\mathbbm{1}$ stands for the indicator function. $Sim(m,n)$ is the function for quantifying the output similarity between layer $m$ and $n$. We choose Centered Kernel Alignment (CKA) [35], a common metric for measuring layer similarities inside or between neural networks [25]. A higher $\kappa$ stands for a more cohesive layer set defined by layers from $i$ to $j$. Inserting AdaNCA after layer $i$ would partition the network into two layer sets, and we can compute the sum of $\kappa$ of the layers before and after $i$, i.e., $\mathcal{K}(i)=\kappa(1,i)+\kappa(i+1,L)$. This serves as a quantification of the network redundancy that corresponds to position $i$. We assume that AdaNCA will not change the layer similarity structure because it is too small compared to a single layer in all ViTs.
In addition to the quantification of the network redundancy, the robustness improvement, brought by AdaNCA, is quantified using the relative increase in the attack failure rate of the AdaNCA-inserted models and the corresponding baseline. Specifically, if a model can achieve $\alpha$ clean test accuracy as well as $\alpha^{\prime}$ accuracy under adversarial attacks, the attack failure rate is $\beta=\frac{\alpha^{\prime}}{\alpha}$. The robustness improvement $\gamma$ is then defined as $\gamma=\frac{\beta_{AdaNCA{}}-\beta_{base}}{\beta_{base}}$. In our experiments, we find that $\gamma$ is significantly correlated with the network redundancy $\mathcal{K}$ (Pearson correlation $r=0.6938,p<0.001$). We refer readers to Appendix A for details of the experiments. The results validate our hypothesis and indicate that AdaNCA should be inserted into the position that can maximize network redundancy. We develop a dynamic programming algorithm to find these positions and refer readers to Appendix A.1 for the details.
@@ 4 Experiments
We use four ViT models as the baseline: Swin-base (Swin-B) [38], FAN-Base-Hybrid (FAN-B) [78], RVT-Base-Plus (RVT-B) [42], and ConViT-Base (ConViT-B) [16]. They include a hierarchical model (Swin), a convolution-attention hybrid model (FAN), and two regular models in which all layers share the same structure (RVT and ConViT). RVT is specifically built to be a robust model, whereas ConViT is not. All four models are equipped with different kinds of local structures. We use two SOTA models in terms of robustness against out-of-distribution (OOD) data, TAPADL-RVT and TAPADL-FAN [23] for comparison. Note that the SOTA method involves training with an additional loss (ADL) and that, for completeness, we keep the results from the TAPADL models trained with such a loss. However, as our focus is on the effect of architectural changes, we do not incorporate the ADL loss in training the AdaNCA-enhanced ViTs. We follow the training scheme for each model, respectively, to train the AdaNCA-equipped model from scratch on ImageNet1K. We conduct the analysis presented in Section 3.3 to decide the optimal position to insert multiple AdaNCA modules, in which the ImageNet1K pre-trained weights for analysis are obtained from the PyTorch Image Models library [69]. To balance between the computational cost and robustness improvement, we limit the number of AdaNCA to two or three, depending on the model architecture. The recurrent time step of all AdaNCA is chosen from $\{[2,2],[2,4],[3,5]\}$, and we follow the design principle of ViT; it is to put more computation in the middle or high layers [38]. We fix the random steps during testing to a single integer chosen from the range to achieve precise results and ensure non-stochasticity during adversarial attacks. All the activation functions used in the MLP in AdaNCA are GELU [29], and the input, hidden, and output dimensionalities of the MLP are all the same. We refer readers to Appendix C.6 for the details of the training and insert scheme of AdaNCA. All of our experiments are performed on four Nvidia A100 GPUs.
Table 1: The performance of AdaNCA-enhanced ViTs and their corresponding baselines. We report the mCE for IM-C (lower is better) and accuracy for other benchmarks. AdaNCA consistently improves the clean accuracy as well as the robustness to adversarial and out-of-distribution inputs of the baseline models. Note that our models also outperform the larger baselines ( sign), indicating that the performance improvement does not merely originate from the increase in the number of parameters or FLOPS. The TAPADL method [23] can lead to more vulnerable models compared to the baselines (green numbers). Bold indicates the best model. : the test is conducted on models pre-trained on ImageNet22K [42], see Appendix C.9.
@@
(M) (G) Clean Acc. PGD [41] CW [5] APGD-DLR [9] APGD-CE [9] IM-A [13] IM-C ($\downarrow$) [28] IM-R [30] IM-SK [67]
@@
Table 2: Comparisons of corruption error on each corruption type of ImageNet-C. AdaNCA consistently improves the performance of the baseline model in all categories (Swin-B), and can achieve better results in most categories compared to the SOTA method (TAPADL-RVT). Bold indicates the best model. Lower is better.
@@ 4.2 Layer similarity structure
Our key assumption in Section 3.3 is that AdaNCA will not change the layer similarity structure due to its small size, and that is why we use pre-trained networks to conduct this analysis. Here, we examine the pair-wise layer similarities in Swin-B [38] and Swin-B-AdaNCA in Figure 5.
Figure 4: Pair-wise layer similarities. Layer sets are marked in red boxes. Swin-B-AdanCA has a clearer stage partition, which might be attributed to AdaNCA acting as an information transmitter between different layer sets.
Figure 5: Ablation on the (a) scales and (b) number of kernels used in our multi-scale Dynamic Interaction. Overly large scales can undermine the performance and so do too many or too few kernels. We choose $\mathcal{S}=2$, $\mathcal{M}=4$ to balance between the clean accuracy and robustness.
@@ 4.3 Ablation studies
We conduct ablation studies on ImageNet100, a 100-class subset of ImageNet1K. Previous studies [7, 15, 72, 74] have shown that ImageNet100 serves as a representative subset of ImageNet1K. Hence, we can obtain representative results for the self-evaluation of the model while efficiently using the computational resources. All the ablation experiments are based on a Swin-tiny [38] model. We insert it after the fourth layer to obtain the best robustness improvement, according to our analysis in Section 3.3 and Appendix A. First, we ablate on two hyperparameters, the number of convolutional kernels used in the Dynamic Interaction stage ($\mathcal{M}$) and the maximum scale ($\mathcal{S}$) used in the multi-scale Dynamic Interaction stage. The Clean Accuracy and Attack Failure Rate are shown in Figure 5. The attack failure rate is quantified using the same method as in Section 3.3 and Appendix A. Multi-scale interaction can contribute to the performance while overly large scale can lose local information and complicate the process of selecting the interaction neighbors. This issue is also observed in a previous work [62]. Increasing the number of kernels benefits the performance while too many kernels undermine the robustness. According to the results, we choose $\mathcal{S}=2$, $\mathcal{M}=4$. We then perform ablations on several design choices:
@@
  • Dynamic interaction (DynIn). Ablate the Dynamic Interaction so that the interaction results are simply summed together. The number of kernels remains the same.
As shown in Table 3, the highest robustness improvement is achieved with all components. Among them, turning off the recurrence leads to the largest drop in the robustness, as recurrence allows the model to explore more cell states than finishing the update in a single step. While it achieves the highest clean accuracy, it uses $\sim$4x more parameters than our method, and the improvement of the clean performance is likely due to more parameters. Without any of the two sources of randomness, stochastic update and random steps, the model cannot adapt to the variability of the inputs and thus exhibits vulnerabilities against adversarial attacks. Finally, turning off our Dynamic Interaction will cause drops in both clean accuracy and robustness, as tokens cannot decide their unique interaction weights and thus cannot generalize to noisy inputs. We provide extended ablation studies in Section D.2 in the Appendix.
Table 3: Ablation studies on design choices. Each of the AdaNCA design choices contributes to improved performance and robustness. Baseline is a Swin-tiny [38] model trained on ImageNet-100. Bold indicates the best model.
Exp. Type Recur StocU RandS DynIn Params (M) FLOPS (G) Accuracy ($\uparrow$) Attack Failure Rate ($\downarrow$)
Baseline 27.59 4.5 86.56 12.29
@@ 4.4 Noise sensitivity examination
A drawback of adversarially robust models is their increased sensitivity to noise in specific frequency bands [60]. For instance, while adversarially trained models are robust against adversarial attacks, they can be sensitive to noise from a larger frequency-band range compared to standard models [60]. Here, we use the method and data from [60] to examine the noise sensitivity of AdaNCA-enhanced ViTs. Specifically, we evaluate the classification performance on a set of images that are mixed with noise of varying magnitudes and frequencies. The noise is sampled from Gaussian distribution with zero mean and the standard deviation indicates its magnitude. Then, it is filtered within various spatial-frequency bands, resulting in different frequencies of noise. Higher classification accuracy on a specific noise type indicates that the model is less sensitive to that noise. Figure 6 presents the results, including human data sourced from [60]. Our findings demonstrate that AdaNCA enhances ViTs by reducing the sensitivity to noise with certain frequency components, equipping ViTs with more human-like noise-resilient abilities. Crucially, AdaNCA improves the model robustness differently than adversarial training since the AdaNCA-enhanced ViTs do not exhibit increased sensitivity to the noise. We quantitatively validate the conclusion and refer readers to Appendix C.16 for the details.
Figure 6: Classification accuracy of humans and ViTs on noisy images. The images are perturbed by Gaussian noise with different standard deviations (Magnitude of Noise) and are filtered with different spatial-frequency bands (Frequency of Noise). AdaNCA improves the accuracy on images with certain types of noise (dotted black boxes), indicating that it makes ViTs less sensitive to them. Quantitative results are presented in Appendix C.16.